# HAIC: Improving Human Action Understanding and Generation with Better Captions for Multi-modal Large Language Models

### v1

@@ 4.2 Results on HAICBench
Caption Evaluation Setting. We follow the caption evaluation setting of HAICBench, where MLLM baselines generate action-related captions for each video, and then the captions and questions are fed into Gemini-1.5-Pro for answers. The results are presented in Table 2. Our post-trained LLaVA-Video-ActionPro-7B achieves state-of-the-art performance among open-source MLLMs, with a 5% performance gain. These results show that training with HAICTrain can significantly improve the caption quality of human actions. Furthermore, using human annotated captions as input can achieve very high accuracy, which means there is still considerable room for improvement. Note that the accuracies of some classes are lower than 25% because we prompted the model not to answer the question if no clues are found in the caption.
Standard Evaluation Setting. In this setting, videos and questions are directly fed into MLLMs for question answering. The results are presented in Table 3. Our post-trained LLaVA-Video-ActionPro-7B outperforms LLaVA-Video by 2.1%, achieving state-of-the-art performance among open-source MLLMs. On the action-related subtasks, our model surpasses the baseline LLaVA-Video-7B model by 1.3%-5.3%. These findings highlight the effectiveness of our HAICTrain dataset for enhancing human action understanding. Additionally, Gemini significantly outperforms GPT-4o, which is the reason why we leverage it to generate the captions of our HAICTrain dataset.
@@ 4.5 Effectiveness for Text-video Generation
To evaluate our method’s effectiveness in text-to-video generation, we use the MovieGenBench dataset Polyak et al. (2024), consisting of 1,003 videos. We first take LLaVA-Video-7B and LLaVA-Video-ActionPro-7B to generate captions, which are then fed to HunyuanVideo Kong et al. and Wanx2.1 Aliyun (2023) to generate videos. Then, five human annotators assess the semantic relevance between videos generated by two captions, classifying them as “Good”, “Same”, or “Bad”. The annotators are asked to make decisions upon the original videos in MovieGen. We measure the GSB score, i.e., the percentage of “Good” and “Same” relative to that of “Bad” and “Same”. Results show that LLaVA-Video-ActionPro-7B outperforms LLaVA-Video-7B in generating captions that lead to more semantically accurate videos. On HunyuanVideo, LLaVA-Video-ActionPro-7B achieves a GSB score of 2.15, and on Wanx2.1, 6.81. LLaVA-Video-7B often produced verbose but less accurate captions, failing to capture the original video’s core actions. These findings highlight our model’s superior ability to comprehend the actions in videos and relay captions that enable high-quality, semantically faithful text-to-video generation. More cases are presented in Appendix E.
@@ 6 Limitations
Appendix A Automatic Annotation
@@
Appendix B Evaluation datasets
@@
Appendix C QA Pair Generation
@@
Appendix D Caption Evaluation Setting
@@
Appendix E Case Study on Text-to-Video Generation
@@
Appendix F Potential Risks

### v2

@@ 4.2 Results on HAICBench
Caption Evaluation Setting. We follow the caption evaluation setting of HAICBench, where MLLM baselines generate action-related captions for each video, and then the captions and questions are fed into Gemini-1.5-Pro for answers. The results are presented in Table 2. Our post-trained LLaVA-Video-ActionPro-7B achieves state-of-the-art performance among open-source MLLMs, with a 6% performance gain. These results show that training with HAICTrain can significantly improve the caption quality of human actions. Furthermore, using human annotated captions as input can achieve very high accuracy, which means there is still considerable room for improvement. Note that the accuracies of some classes are lower than 25% because we prompted the model not to answer the question if no clues are found in the caption.
Standard Evaluation Setting. In this setting, videos and questions are directly fed into MLLMs for question answering. The results are presented in Table 3. Our post-trained LLaVA-Video-ActionPro-7B outperforms LLaVA-Video by 2.2%, achieving state-of-the-art performance among open-source MLLMs. On the action-related subtasks, our model surpasses the baseline LLaVA-Video-7B model by 1%-2%. These findings highlight the effectiveness of our HAICTrain dataset for enhancing human action understanding. Additionally, Gemini significantly outperforms GPT-4o, which is the reason why we leverage it to generate the captions of our HAICTrain dataset.
@@ 4.5 Effectiveness for Text-video Generation
To evaluate our method’s effectiveness in text-to-video generation, we use the MovieGenBench dataset Polyak et al. (2024), consisting of 1,003 videos. We first take LLaVA-Video-7B and LLaVA-Video-ActionPro-7B to generate captions, which are then fed to HunyuanVideo Kong et al. and Wanx2.1 Aliyun (2025) to generate videos. Then, five human annotators assess the semantic relevance between videos generated by two captions, classifying them as “Good”, “Same”, or “Bad”. The annotators are asked to make decisions upon the original videos in MovieGen. We measure the GSB score, i.e., the percentage of “Good” and “Same” relative to that of “Bad” and “Same”. Results show that LLaVA-Video-ActionPro-7B outperforms LLaVA-Video-7B in generating captions that lead to more semantically accurate videos. On HunyuanVideo, LLaVA-Video-ActionPro-7B achieves a GSB score of 2.15, and on Wanx2.1, 6.81. LLaVA-Video-7B often produced verbose but less accurate captions, failing to capture the original video’s core actions. These findings highlight our model’s superior ability to comprehend the actions in videos and relay captions that enable high-quality, semantically faithful text-to-video generation. More cases are presented in Appendix F.
@@ 6 Limitations
7 Acknowledgment
This work was supported in part by the National Natural Science Foundation of China(Grant Nos. 62376069 and 62236003), in part by the Young Elite Scientists Sponsorship Program by CAST (Grant No. 2023QNRC001), in part by Guangdong Basic and Applied Basic Research Foundation (Grant No. 2024A1515012027), in part by Jiangsu Science and Technology Major Program (Grant No. BG2024041), and in part by the Shenzhen Science and Technology Program (Grant Nos. KQTD20240729102207002 and ZDSYS20230626091203008).
Appendix A Comparison with concurrent benchmarks
We compared our benchmark with three concurrent works on human action understanding—MoVidBench Chen et al. (2024c), MotionBench Hong et al. (2025), and FAVORBench Tu et al. (2025)—across three key aspects of human actions:
  • Action details: Fine-grained motion annotations (e.g., body movements, intensity, direction).
  • Action transitions: Details on the changes of two actions.
  • Multi-human interactions: Actions involving multiple humans.
Our benchmark is the only one covering all three aspects, as illustrated in Table 5 below.
Benchmark Action  Details Action  Transitions Multi-Human  Interactions
MoVidBench $\checkmark$ $\checkmark$ None
MotionBench $\times$ $\checkmark$ Few
FAVORBench $\checkmark$ $\checkmark$ Few
HAICBench $\checkmark$ $\checkmark$ Rich
Table 5: Comparison of recent human action understanding benchmarks on three aspects of human action
Appendix B Automatic Annotation
@@
Appendix C Evaluation datasets
@@
Appendix D QA Pair Generation
@@
Appendix E Caption Evaluation Setting
@@
Appendix F Case Study on Text-to-Video Generation
@@
Appendix G Potential Risks